# Separating complexity classes of LCL problems on grids

### v1

@@ Abstract.
KB’s research is partially supported by the NSF Graduate Research Fellowship Program under grant DGE-2146752. AB’s research is partially supported by the NSF CAREER grant DMS-2239187. FW is supported by the NSF under award number DMS-2402064.
$\mathtt{CONTINUOUS}(\mathbb{Z}^{n})$$\mathtt{BOREL}(\mathbb{Z}^{n})$$\mathtt{MEASURE}(\mathbb{Z}^{n})$$\mathtt{BAIRE}(\mathbb{Z}^{n})$$\mathtt{FIID}(\mathbb{Z}^{n})$$\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$$\mathtt{FFIID}(\mathbb{Z}^{n})$$\mathtt{COMPUTABLE}(\mathbb{Z}^{n})$Thm. 1.25Thm. 1.8Thm. 1.18Thm. 1.8$=$?$=$?Thm. 1.18Thm. 1.8Thm. 1.14Thms. 1.19 and 1.25[Sequel]Thm. 1.26Thm. 1.26Thm. 1.23Thm. 1.26 and [Sequel]Thm. 1.26 and [Sequel]$A$$B$means $A\subsetneq B$$A$$B$means $A\subseteq B$$A$$B$means $A\not\subseteq B$
Figure 1. Complexity classes of LCL problems on $\mathbb{Z}^{n}$ with $n\geqslant 2$. All known inclusions/non-inclusions between these classes follow from the ones shown. The relations established in this paper and in the forthcoming paper [Sequel] are labeled accordingly.
@@ 1.1.  Descriptive complexity of LCL problems
    We stress that in the definition of $\mathtt{MEASURE}(\Gamma)$, the action $\Gamma\mathrel{\reflectbox{$\righttoleftarrow$}}(X,\mu)$ need not be measure/preserving. (However, see Question 6.11.)
@@ 1.2.  Shift actions: the universal domain for descriptive combinatorics?
Although the above definition makes sense in general, we will only use it for countable (in fact, finitely generated) groups $\Gamma$. Note that if $\Gamma$ is countable and $X$ is a standard Borel, (zero-dimensional) Polish, or standard probability space, then so is $\mathcal{S}(X,\Gamma)$ equipped with a suitable product structure, and the shift action $\Gamma\mathrel{\reflectbox{$\righttoleftarrow$}}\mathcal{S}(X,\Gamma)$ is Borel (as well as continuous and measure-preserving when these make sense).
@@
For every integer $n\geqslant 2$, we have $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})\subsetneq\mathtt{BAIRE}(\mathbb{Z}^{n})$.
@@
As for the relationship between $\mathtt{FIID}(\Gamma)$ and $\mathtt{FFIID}(\Gamma)$, prior to this work, the following has been an open question:
@@ 1.4.  Computable combinatorics
It is consistent with all the previously mentioned results that the class $\mathtt{COMPUTABLE}(\mathbb{Z}^{n})$ includes all the complexity classes defined in this paper. For comparison, note that $\mathtt{COMPUTABLE}(\mathbb{F}_{n})=\mathtt{BAIRE}(\mathbb{F}_{n})$ does include all the other complexity classes for the group $\mathbb{F}_{n}$. Nevertheless, in the forthcoming second part of this study [Sequel], we will show that $\mathtt{COMPUTABLE}(\mathbb{Z}^{n})$ includes neither $\mathtt{MEASURE}(\mathbb{Z}^{n})$ nor $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$:
Theorem 1.24 ([Sequel]).
For every integer $n\geqslant 2$, we have
$\mathtt{MEASURE}(\mathbb{Z}^{n})\setminus\mathtt{COMPUTABLE}(\mathbb{Z}^{n})\,\neq\,\varnothing\qquad\text{and}\qquad\mathtt{BaireSHIFT}(\mathbb{Z}^{n})\setminus\mathtt{COMPUTABLE}(\mathbb{Z}^{n})\,\neq\,\varnothing.$
The proof of Theorem 1.24 builds on the techniques introduced in this paper but requires a considerable amount of additional machinery to be developed; that is why we decided to present it in a separate paper.
@@ 1.5.  Summary, further results, and an overview of the remainder of the paper
All the results mentioned above, with the exception of Theorems 1.19 and 1.24, follow from the existence of a single LCL problem $\Pi$:
@@
It is clear that Theorem 1.25 implies Theorems 1.8, 1.14, 1.18, and 1.23. By modifying the problem witnessing Theorem 1.25, we obtain the following additional results:
@@
In particular, this proves that $\mathtt{COMPUTABLE}(\mathbb{Z}^{n})\not\subseteq\mathtt{FIID}(\mathbb{Z}^{n})$, and thus certainly
$\mathtt{COMPUTABLE}(\mathbb{Z}^{n})\,\not\subseteq\,\mathtt{MEASURE}(\mathbb{Z}^{n}).$
Together with Theorem 1.24, this shows that the classes $\mathtt{MEASURE}(\mathbb{Z}^{n})$ and $\mathtt{COMPUTABLE}(\mathbb{Z}^{n})$ are incomparable (by contrast, for the free group $\mathbb{F}_{n}$, the complexity classes defined in this paper—with the possible exception of $\mathtt{FFIID}(\mathbb{F}_{n})$—are linearly ordered by inclusion). Theorem 1.26 also makes a first step toward refuting the possibility that $\mathtt{BAIRE}(\mathbb{Z}^{n})$ is a subset of $\mathtt{MEASURE}(\mathbb{Z}^{n})$, since it yields an LCL problem that is in $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$ but not in $\mathtt{MEASURE}(\mathbb{Z}^{n})$.
The main challenge in proving Theorems 1.25 and 1.26 is to achieve the desired behavior with locally checkable problems. For example, a version of Theorem 1.14 (non-universality of the shift in the Baire-measurable context) for non-local problems was proved before in [BerGeneric, Corl. 2.12]; however, that result holds for arbitrary countably infinite groups, which shows that the subtle distinctions between, say, $\mathbb{Z}^{n}$ and $\mathbb{F}_{n}$ are lost when non-local problems are considered.
The starting point for our constructions is a type of combinatorial structure called a toast. This concept originates in the work of Conley and Miller [CM16] (the term “toast” was coined by Miller) and has been widely used in descriptive combinatorics ever since [bowen2021perfect, BPZ, trees, ContinuousAbelian, BorelAbelian, gao2022forcing, marks2017borel]. Related notions also appear in the theory of random processes [FIIDSchreier, Holroyd, Spinka]. Recently, Grebík and Rozhoň initiated a systematic study of toast-based constructions, under the name of TOAST algorithms [GRgrids], and closely related ASI algorithms were studied by Qian and the fourth named author [ASIalgorithms]. A central role in our arguments is played by toasts with rectangular pieces, which we call rectangular toasts for brevity. We define rectangular toasts and establish their relevant properties in §2. Then, in §3, we use them to design an LCL problem witnessing Theorem 1.25. In §4 we explain how to modify the construction in the proof of Theorem 1.25 to derive Theorem 1.26. We conclude the paper with the proof of Theorem 1.19 in §5 and then some final remarks, including a list of open problems, in §6.
@@ 2.2.  Rectangular toasts and measure theory
$\mathcal{Q}_{i}\,\coloneqq\,\big{\{}C\in\bigcup_{j<i}\mathcal{D}\,:\,\text{$\mathsf{dist}(\partial C,\partial D)\leqslant q$ for some $D\in\mathcal{D}_{i}$}\big{\}}\quad\text{and}\quad\mathcal{Q}\,\coloneqq\,\bigcup_{i\in\mathbb{N}}\mathcal{Q}_{i}.$
@@ 3.1.  The rectangular $q$-toast problem
Options (RT1) and (RT2) are mutually exclusive since the rectangles in a rectangular $q$-toast are finite. Let $\mathcal{C}$ be the set of all components of the graph $G_{n}[f^{-1}(\mathtt{R})]$. For each $C\in\mathcal{C}$, fix a set $S_{C}$ such that $\partial S_{C}=V(C)$ and $\partial^{+}S_{C}$ is the set of all blue points with a neighbor in $V(C)$, given by Lemma 3.3. Suppose for some $C\in\mathcal{C}$, $S\coloneqq S_{C}$ is infinite. Since $S$ is a product of intervals at least one of which is infinite and $\partial S\neq\varnothing$, we can choose $x\in\partial S$ and $\mathsf{d}\in\{\pm\mathsf{e}_{1},\ldots,\pm\mathsf{e}_{n}\}$ so that the set $L\coloneqq\{i\mathsf{d}+x\,:\,i\in\mathbb{N}\}$ is contained in $\partial S\subseteq f^{-1}(\mathtt{R})$. Furthermore, there is some $\mathsf{d}^{\prime}\in\{\pm\mathsf{e}_{1},\ldots,\pm\mathsf{e}_{n}\}$ orthogonal to $\mathsf{d}$ such that $L+\mathsf{d}^{\prime}\subseteq\partial^{+}S\subseteq f^{-1}(\mathtt{B})$. The definition of $\mathsf{RT}(q)$ implies that for any $y\in\mathbb{Z}^{n}$, if $f(y)=\mathtt{R}$ and $f(\mathsf{d}^{\prime}+y)=\mathtt{B}$, then $f(2\mathsf{d}^{\prime}+y)=\mathtt{G}$. Hence, in this case (RT2) holds.
@@ 3.3.  The colored rectangular $q$-toast problem
4.  Proof of Theorem 1.26
In this section we prove Theorem 1.26. For simplicity, we only present the proof in the case $n=2$; the general argument is the same, mutatis mutandis. In our construction we start with the problem $\mathsf{CRT}(q)$ used to witness Theorem 1.25 and then add some extra structure to it in order to ensure that the resulting problem can no longer be solved measurably on the free part of the shift, but still has a Baire-measurable solution on the free part of the shift as well as a computable solution on any free computable action.
@@
Note that for almost all $x\in E$ and for all $1\leqslant i\leqslant k$, we have $\pi(x)\in U_{\gamma_{i}\cdot\varphi_{i}}$, or, equivalently,
$\pi(x)\,\in\,U_{(\gamma_{1}\cdot\varphi_{1})\cup\ldots\cup(\gamma_{k}\cdot\varphi_{k})}.$
Therefore, $(\gamma_{1}\cdot\varphi_{1})\cup\ldots\cup(\gamma_{k}\cdot\varphi_{k})\in\mathcal{I}$, and the proof is complete. ∎
@@ 6.2.  Open problems
Note that $\mathtt{CONTINUOUS}(\mathbb{Z}^{n})$ is the minimum element of the poset of inclusions shown in Fig. 1. Currently, it appears possible that $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$ is the maximum element of this poset. Is it? To wit, do we have $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})\supset\mathtt{COMPUTABLE}(\mathbb{Z}^{n})$ and $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})\supset\mathtt{FIID}(\mathbb{Z}^{n})$?
Question 6.8.
@@
There are also general questions that are of interest for all finitely generated groups. Here are two examples, the first of which is closely related to Question 6.4.
@@
Question 6.11.
Let $\mathtt{MEASURE}_{\mathsf{pmp}}(\Gamma)$ be the class of all LCL problems on $\Gamma$ that admit a measurable solution on every free measure-preserving action $\Gamma\mathrel{\reflectbox{$\righttoleftarrow$}}(X,\mu)$, where $(X,\mu)$ is a standard probability space. Clearly, $\mathtt{MEASURE}(\Gamma)\subseteq\mathtt{MEASURE}_{\mathsf{pmp}}(\Gamma)$. Can this inclusion be strict?

### v4

@@ Abstract.
KB’s research is partially supported by the NSF Graduate Research Fellowship Program under grant DGE-2146752. AB’s research is partially supported by the NSF CAREER grant DMS-2528522. CL is supported by the ERC Advanced Grant “ERMiD”. FW is supported by the NSF under award number DMS-2402064.
@@ 1.1.  Descriptive complexity of LCL problems
    We stress that in the definition of $\mathtt{MEASURE}(\Gamma)$, the action $\Gamma\mathrel{\reflectbox{$\righttoleftarrow$}}(X,\mu)$ need not be measure/preserving. (However, see Question 6.10.)
@@ 1.2.  Shift actions: the universal domain for descriptive combinatorics?
We will only use the above definition for countable (in fact, finitely generated) groups $\Gamma$. If $\Gamma$ is countable and $X$ is a standard Borel, (zero-dimensional) Polish, or standard probability space, then so is $\mathcal{S}(X,\Gamma)$ equipped with a suitable product structure, and the shift action $\Gamma\mathrel{\reflectbox{$\righttoleftarrow$}}\mathcal{S}(X,\Gamma)$ is Borel (as well as continuous and measure-preserving when these make sense).
@@
For every integer $n\geqslant 2$, we have $\mathtt{BAIRE}(\mathbb{Z}^{n})\subsetneq\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$.
@@
[…]
@@ 1.5.  Summary, further results, and an overview of the remainder of the paper
All the results mentioned above, with the exception of Theorem 1.19, follow from the existence of a single LCL problem $\Pi$:
@@
It is clear that Theorem 1.24 implies Theorems 1.8, 1.14, 1.18, and 1.23. The main challenge in proving Theorem 1.24 is to achieve the desired behavior with a locally checkable problem. For example, a version of Theorem 1.14 (non-universality of the shift in the Baire-measurable context) for non-local problems was proved before in [BerGeneric, Corl. 2.12]; however, that result holds for arbitrary countably infinite groups, which shows that the subtle distinctions between, say, $\mathbb{Z}^{n}$ and $\mathbb{F}_{n}$ are lost when non-local problems are considered.
The starting point for our construction is a type of combinatorial structure called a toast. This concept originates in the work of Conley and Miller [CM16] (the term “toast” was coined by Miller) and has been widely used in descriptive combinatorics ever since [bowen2021perfect, BPZ, trees, ContinuousAbelian, BorelAbelian, gao2022forcing, marks2017borel]. Related notions also appear in the theory of random processes [FIIDSchreier, Holroyd, Spinka]. Recently, Grebík and Rozhoň initiated a systematic study of toast-based constructions, under the name of TOAST algorithms [GRgrids], and closely related ASI algorithms were studied by Qian and the fourth named author [ASIalgorithms]. A central role in our arguments is played by toasts with rectangular pieces, which we simply call rectangular toasts. Briefly, a rectangle in a free action $\mathbb{Z}^{n}\mathrel{\reflectbox{$\righttoleftarrow$}}X$ is a set of the form $R\cdot x$, where $x\in X$ and $R\subset\mathbb{Z}^{n}$ is an $n$-dimensional axis-parallel rectangle in $\mathbb{Z}^{n}$, and a rectangular toast $\mathcal{R}$ for the action $\mathbb{Z}^{n}\mathrel{\reflectbox{$\righttoleftarrow$}}X$ is a nested family of rectangles (see Fig. 3). We formally define rectangular toasts and establish their relevant properties in §2. It turns out that one can always construct Borel rectangular toasts $\mathcal{R}$ such that the union of the rectangles in $\mathcal{R}$ is conull (Proposition 2.6), but it may be impossible to make it comeager (Proposition 2.12). In other words, building a rectangular toast with a large union is a problem that can be solved measurably but not Baire-measurably, indicating that rectangular toasts may be employed toward proving Theorem 1.8. Unfortunately, the definition of a rectangular toast is plainly non-local. Nevertheless, in §3, we find a way to incorporate rectangular toasts in an LCL problem and show that the resulting problem witnesses Theorem 1.24.
$\mathtt{CONTINUOUS}(\mathbb{Z}^{n})$$\mathtt{BOREL}(\mathbb{Z}^{n})$$\mathtt{MEASURE}(\mathbb{Z}^{n})$$\mathtt{BAIRE}(\mathbb{Z}^{n})$$\mathtt{FIID}(\mathbb{Z}^{n})$$\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$$\mathtt{FFIID}(\mathbb{Z}^{n})$$\mathtt{COMPUTABLE}(\mathbb{Z}^{n})$Thm. 1.24Thm. 1.8Thm. 1.18Thm. 1.8$=$?$=$?Thm. 1.18Thm. 1.8Thm. 1.14[Sequel]Thm. 1.25 and [Sequel]Thm. 1.25 and [Sequel]Thm. 1.23Thm. 1.25 and [Sequel]Thms. 1.19 and 1.24Thm. 1.25 and [Sequel]
Figure 1. Complexity classes of LCL problems on $\mathbb{Z}^{n}$ with $n\geqslant 2$. Here:
  • $A$$B$
    indicates that $A$ is a strict subset of $B$,
  • $A$$B$$=$?
    means that $A\subseteq B$ and we do not know whether $A=B$,
  • $A$$B$
    indicates that $A$ is not a subset of $B$.
All known inclusions/non-inclusions between these classes follow from the ones shown. The relations established in this paper and in the forthcoming paper [Sequel] are labeled accordingly.
The LCL problem we use to prove Theorem 1.24 can be enhanced with extra layers of structure to derive further separation results between different complexity classes. In particular, in §4 we modify the problem witnessing Theorem 1.24 to obtain the following:
@@
It follows that we have $\mathtt{COMPUTABLE}(\mathbb{Z}^{n})\not\subseteq\mathtt{FIID}(\mathbb{Z}^{n})$, and hence $\mathtt{COMPUTABLE}(\mathbb{Z}^{n})\not\subseteq\mathtt{MEASURE}(\mathbb{Z}^{n})$. Theorem 1.25 also makes a first step toward refuting the possibility that $\mathtt{BAIRE}(\mathbb{Z}^{n})$ is a subset of $\mathtt{MEASURE}(\mathbb{Z}^{n})$, since it yields an LCL problem that is in $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$ but not in $\mathtt{MEASURE}(\mathbb{Z}^{n})$.
We conclude the paper with the proof of Theorem 1.19 in §5 and then some final remarks, including a list of open problems, in §6.
The method of adding extra structure to the LCL problem from Theorem 1.24 that we use to establish Theorem 1.25 is very flexible and can be employed to obtain a variety of other results. Specifically, in the forthcoming second part of this study [Sequel], we will prove the following:
Theorem 1.26 ([Sequel]).
For every integer $n\geqslant 2$, the classes $\mathtt{MEASURE}(\mathbb{Z}^{n})$, $\mathtt{COMPUTABLE}(\mathbb{Z}^{n})$, and $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$ are pairwise incomparable with respect to inclusion.
Although the proof of Theorem 1.26 builds on the work in this paper, it requires a considerable amount of additional technical machinery to be developed; that is why we decided to present it separately. Theorem 1.26 clearly strengthens Theorem 1.8 (since $\mathtt{BAIRE}(\mathbb{Z}^{n})$ is a strict subset of $\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$). In view of the inclusion $\mathtt{FFIID}(\mathbb{Z}^{n})\subset\mathtt{BaireSHIFT}(\mathbb{Z}^{n})$ (Theorem 1.19), it also strengthens Theorem 1.18. The incomparability between the three classes in Theorem 1.26 makes the contrast between the groups $\mathbb{Z}^{n}$ and $\mathbb{F}_{n}$ especially stark, since for the free group $\mathbb{F}_{n}$, the complexity classes defined in this paper—with the possible exception of $\mathtt{FFIID}(\mathbb{F}_{n})$—are linearly ordered by inclusion. Theorem 1.26 also shows that, while $\mathtt{CONTINUOUS}(\mathbb{Z}^{n})$ is the minimum element in the poset of the complexity classes studied in this paper, it does not have a maximum element. (By contrast, for $\mathbb{F}_{n}$, the class $\mathtt{BAIRE}(\mathbb{F}_{n})=\mathtt{BaireSHIFT}(\mathbb{F}_{n})=\mathtt{COMPUTABLE}(\mathbb{F}_{n})$ is maximum.)
The current state of our knowledge regarding the complexity classes of LCL problems on $\mathbb{Z}^{n}$ for $n\geqslant 2$, including the results of this paper and those of [Sequel], is summarized in Fig. 1.
@@ 2.2.  Rectangular toasts and measure theory
$\mathcal{Q}_{i}\,\coloneqq\,\big{\{}C\in\bigcup_{j<i}\mathcal{D}_{j}\,:\,\text{$\mathsf{dist}(\partial C,\partial D)\leqslant q$ for some $D\in\mathcal{D}_{i}$}\big{\}}\quad\text{and}\quad\mathcal{Q}\,\coloneqq\,\bigcup_{i\in\mathbb{N}}\mathcal{Q}_{i}.$
@@ 3.1.  The rectangular $q$-toast problem
Options (RT1) and (RT2) are mutually exclusive since the rectangles in a rectangular $q$-toast are finite. Let $\mathcal{C}$ be the set of all components of the graph $G_{n}[f^{-1}(\mathtt{R})]$. For each $C\in\mathcal{C}$, fix a set $S_{C}$ such that $\partial S_{C}=V(C)$ and $\partial^{+}S_{C}$ is the set of all blue points with a neighbor in $V(C)$, given by Lemma 3.3. Suppose for some $C\in\mathcal{C}$, $S\coloneqq S_{C}$ is not a (finite) rectangle. Since $S$ is a product of intervals at least one of which is infinite and $\partial S\neq\varnothing$, we can choose $x\in\partial S$ and $\mathsf{d}\in\{\pm\mathsf{e}_{1},\ldots,\pm\mathsf{e}_{n}\}$ so that the set $L\coloneqq\{i\mathsf{d}+x\,:\,i\in\mathbb{N}\}$ is contained in $\partial S\subseteq f^{-1}(\mathtt{R})$. Furthermore, there is some $\mathsf{d}^{\prime}\in\{\pm\mathsf{e}_{1},\ldots,\pm\mathsf{e}_{n}\}$ orthogonal to $\mathsf{d}$ such that $L+\mathsf{d}^{\prime}\subseteq\partial^{+}S\subseteq f^{-1}(\mathtt{B})$. The definition of $\mathsf{RT}(q)$ implies that for any $y\in\mathbb{Z}^{n}$, if $f(y)=\mathtt{R}$ and $f(\mathsf{d}^{\prime}+y)=\mathtt{B}$, then $f(2\mathsf{d}^{\prime}+y)=\mathtt{G}$. Hence, in this case (RT2) holds.
@@ 3.3.  The colored rectangular $q$-toast problem
4.  Adding extra structure: Proof of Theorem 1.25
In this section we prove Theorem 1.25. For simplicity, we only present the proof in the case $n=2$; the general argument is the same, mutatis mutandis. In our construction we start with the problem $\mathsf{CRT}(q)$ used to witness Theorem 1.24 and then add some extra structure to it in order to ensure that the resulting problem can no longer be solved measurably on the free part of the shift, but still has a Baire-measurable solution on the free part of the shift as well as a computable solution on any free computable action. Similar (but much more complicated) arguments will be used in [Sequel] to prove Theorem 1.26.
@@
Note that for almost all $x\in E$ and for all $1\leqslant i\leqslant k$, we have $\pi(x)\in U_{\gamma_{i}\cdot\varphi_{i}}$, or, equivalently, $\pi(x)\in U_{(\gamma_{1}\cdot\varphi_{1})\cup\ldots\cup(\gamma_{k}\cdot\varphi_{k})}$. Therefore, $(\gamma_{1}\cdot\varphi_{1})\cup\ldots\cup(\gamma_{k}\cdot\varphi_{k})\in\mathcal{I}$, and the proof is complete. ∎
@@ 6.2.  Open problems
Question 6.8.
Furthermore, there remain general questions that are of interest for all finitely generated groups. Here are two examples, the first of which is closely related to Question 6.4.
@@
Let $\mathtt{MEASURE}_{\mathsf{pmp}}(\Gamma)$ be the class of all LCL problems on $\Gamma$ that admit a measurable solution on every free measure-preserving action $\Gamma\mathrel{\reflectbox{$\righttoleftarrow$}}(X,\mu)$, where $(X,\mu)$ is a standard probability space. Clearly, $\mathtt{MEASURE}(\Gamma)\subseteq\mathtt{MEASURE}_{\mathsf{pmp}}(\Gamma)$. Can this inclusion be strict? Once again, we know $\mathtt{MEASURE}_{\mathsf{pmp}}(\mathbb{Z})=\mathtt{MEASURE}(\mathbb{Z})$ because $\mathtt{FIID}(\mathbb{Z})=\mathtt{MEASURE}(\mathbb{Z})$, but we do not know the answer for any group that is not virtually $\mathbb{Z}$.